# Topology-Aware Conformal Prediction for Stream Networks

### v1

@@ A.3 Tailup model
[…]
Due to the independence of increments for Brownian motion, only when $r=s\in\wedge u\cap\wedge v$, the covariance is non-zero. Note that for Brownian motion, we have $\text{Cov}(dB(r),dB(s))=1_{r=s}drds$. Hence Lemma A.4 follows. ∎
Lemma A.5.
If we set the moving function $m(r-u)\;=\;\beta\,\exp\!\Bigl{(}-\frac{d(r,u)}{\phi}\Bigr{)},$ with parameters $\beta>0$ (a scale factor) and $\phi>0$ (a range or decay parameter), then the covariance matrix between two locations $u,v$ can be expressed as Equation 4.4.
@@
By Lemma A.4 and substitute $m(r-u)=\beta\,e^{-d(r,u)/\phi}$ and $m(r-v)=\beta\,e^{-d(r,v)/\phi}$, we have
[…]
@@
Appendix B Taxonomy for Related Works
1. [35] 3. [7] 4. [23] 5. [10, 41, 40] 6. [39] 7. Our method, STACI
Figure 7 provides an overview of the conformal prediction literature, as a supplement to the related work. The Venn graph categorizes existing CP methods based on their applicability to different data types and the assumptions they rely on. Specifically, it distinguishes between methods designed for time series data and multivariate data, and further classifies them based on whether they assume no temporal distribution shift in time series or no spatial dependencies in multivariate data.
Traditional CP methods typically require that time series data exhibit no temporal distribution shift or that multivariate data lack spatial dependencies to ensure the exchangeability assumption. However, tailored for spatio-temporal stream networks, our proposed CP method lies at the intersection of these categories in the Venn diagram. Unlike conventional approaches, our method explicitly accounts for both spatial dependencies and temporal shifts, leveraging the underlying topological structure of the network to enhance predictive performance.
Appendix C Additional Experiment Details and Results
C.1 Simulation Data Generation Details
Segment $r_{1}$ and $r_{2}$ starts with (0, 1) and (0.5, 0.8), respectively, and both end with (0.3, 0.5). The next segment $r_{3}$ also starts with (0.3, 0.5), and end with (0.2, 0.1). Segment $r_{4}$ start from (0.6, 0.6), and ends at the same location as $r_{3}$. Starting from this location, $r_{5}$ ends at (0.4, 0). The weights for segment $1-5$ are set as $0.35,0.5,0.85,0.15$ and $1$, respectively. Each segment has two observation locations – one at the start point, another at the middle point.
@@ C.1 Simulation Data Generation Details
C.2 Additional Ablation Study: Offline Experiment
@@
From Figure 5, MultiDimSPCI achieve the closest to our proposed method, STACI, in efficiency. Therefore, we focus our comparison on four specific variants: vanilla MultiDimSPCI($\gamma=0$), MultiDimSPCI($\gamma=0.01$), STACI($\gamma=0$), and STACI($\gamma=0.01$). In the offline setting, STACI does not update the covariance matrix estimation. To ensure a fair comparison, we similarly fix the covariance matrix for MultiDimSPCI methods at the beginning of the test phase.

### v2

@@ A.3 Tail-up model
$\Sigma(u,v)=\text{Cov}(\int_{\wedge u}m(s-u)\sqrt{\frac{w(s)}{w(u)}}dB(s),\int_{\wedge v}m(r-v)\sqrt{\frac{w(r)}{w(v)}}dB(r))$
Due to the independence of increments for Brownian motion, only when $r=s\in\wedge u\cap\wedge v$, the covariance is non-zero. Note that for Brownian motion, we have $\text{Cov}(dB(r),dB(s))=1_{r=s}drds$. Hence Lemma 5 follows. ∎
Lemma 6.
If we set the moving function $m(r-u)\;=\;\beta\,\exp\!\Bigl(-\frac{d(r,u)}{\phi}\Bigr),$ with parameters $\beta>0$ (a scale factor) and $\phi>0$ (a range or decay parameter), then the covariance matrix between two locations $u,v$ can be expressed as:
$\Sigma(u,v)=\sigma^{2}\sqrt{\frac{w(u)}{w(v)}}\exp\bigl(-d(u,v)/\phi\bigr)\,\mathbf{1}_{\{u\to v\}}.$ (29)
@@
By Lemma 5 and substitute $m(r-u)=\beta\,e^{-d(r,u)/\phi}$ and $m(r-v)=\beta\,e^{-d(r,v)/\phi}$, we have
$\Sigma(u,v)\;=\;\int_{\wedge u\,\cap\,\wedge v}\beta^{2}\;\exp\Bigl(-\tfrac{d(r,u)}{\phi}\Bigr)\;\exp\Bigl(-\tfrac{d(r,v)}{\phi}\Bigr)\;\frac{w(r)}{\sqrt{w(u)\,w(v)}}\;dr.$
@@
    $\Sigma(u,v)\;=\;\beta^{2}\,\sqrt{\frac{w(u)}{w(v)}}\,\exp\!\Bigl(-\frac{d(u,v)}{\phi}\Bigr)\int_{\wedge u}\exp\Bigl(-\tfrac{2d(r,u)}{\phi}\Bigr)\;\frac{w(r)}{w(u)}\;dr,$
Note that $\int_{\wedge u}\exp\Bigl(-\tfrac{2d(r,u)}{\phi}\Bigr)\;\frac{w(r)}{w(u)}\;dr=\Sigma(u,u)$ is a constant, since the additivity constraint on $w(u)$ assures the constant variance of site $u$. Thus, the tail‐up exponential model yields a covariance of the form
Appendix B Taxonomy for Related Works
1. [vovk2005algorithmic] 3. [diquigiovanni2021distribution] 4. [messoudi2021copula] 5. [gibbs2021adaptive, zaffran2022adaptive, xu2023conformal, angelopoulos2023conformal, yang2024bellman] 6. [xuconformal] 7. Our method, STACI
Figure 4 provides an overview of the conformal prediction (CP) literature, complementing the discussion in the related work section. The Venn diagram categorizes existing CP methods based on the type of data they are designed for—time series, multivariate data, or both—and the nature of the assumptions they make, particularly regarding exchangeability, temporal stationarity, and spatial independence.
Traditional CP methods, such as split conformal prediction for i.i.d. regression[vovk2005algorithmic], assume full exchangeability, and thus cannot be directly applied to time series or spatially structured data without modification. In the time series domain, recent works such as [angelopoulos2023conformal, zaffran2022adaptive, gibbs2021adaptive, yang2024bellman] relax the exchangeability assumption via online calibration, sliding window methods, scorecasters or dynamic programming. These methods handle distribution shift over time but typically operate in univariate or low-dimensional settings. In contrast, CP methods for multivariate or high-dimensional data [messoudi2021copula, diquigiovanni2021distribution] often focus on constructing prediction sets that exploit geometry or sparsity, but generally assume no temporal structure.
Our proposed method, STACI, is positioned at the intersection of these axes in the Venn diagram. It is designed for high-dimensional time-indexed multivariate data arising from spatio-temporal stream networks. Our method explicitly accounts for both spatial dependencies and temporal shifts, leveraging the underlying topological structure of the network to enhance predictive performance.
Appendix C Adaptive Uncertainty Set Construction
Algorithm 1 STACI
Output: Prediction set $\mathcal{C}(X_{T+1};\alpha)$.
1: // Training
[…]
3: // Calibration
@@ Appendix C Adaptive Uncertainty Set Construction
[…]
10: // Testing
[…]
We present the analysis of the average coverage guarantee of STACI without any assumption about $\epsilon_{t}$. The proof follows from Proposition 4.1 in [gibbs2021adaptive].
Proposition 1.
$\lim_{n^{\prime}\to\infty}\sum_{t\in\mathcal{D}_{\text{test}}}\mathds{1}\{Y_{t}\notin\mathcal{C}(X_{t};\alpha_{t})\}\big/n^{\prime}=1-\alpha.$
While Proposition 1 provides a weaker coverage guarantee compared to Theorem 1, it offers broader applicability, remaining valid even in adversarial online settings. Empirical results suggest that when the error process exhibits minimal distribution shift and the assumptions of Theorems 2 and 1 are only slightly violated, STACI maintains the predefined coverage level ($\gamma=0.01$) while achieving efficient prediction sets. However, when $\gamma>0$, Proposition 1 does not ensure a finite-sample coverage gap. Understanding this limitation and developing methods to control the finite-sample coverage gap presents an interesting direction for future research.
Appendix D Additional Experiment Details and Results
D.1 Computational Resource
Experiments were conducted on a single NVIDIA GeForce RTX 4080 Super GPU, an AMD Ryzen 9 7950X 16-Core Processor CPU, 64GB Memory and 2TB SSD.
D.2 Datasets and Codes
The PeMS03 dataset used in this paper is collectd by California Transportation Agencies (CalTrans) Performance Measurement System (PeMS). It contains three months of statistics on traffic flow every 5 minutes ranging from Sept. 1st 2018 to Nov. 30th 2018, including 358 sensors. The datasets are from https://github.com/guoshnBJTU/ASTGNN/tree/main without available license.
The first backbone ST-Graph model is AGCRN [bai2020adaptive], which we adapted the official implementation from https://github.com/LeiBAI/AGCRN under MIT license. ASTGCN [Guo_Lin_Feng_Song_Wan_2019] was implemented with PyTorch Geometric Temporal package [rozemberczki2021pytorch] from https://github.com/benedekrozemberczki/pytorch_geometric_temporal under MIT license. STGODE [fang2021spatial] was adapted from the official implementation https://github.com/square-coder/STGODE under Apache-2.0 license.
D.3 Simulation Data Generation Details
Segment $r_{1}$ and $r_{2}$ starts with (0, 1) and (0.5, 0.8), respectively, and both end with (0.3, 0.5). The next segment $r_{3}$ also starts with (0.3, 0.5), and end with (0.2, 0.1). Segment $r_{4}$ start from (0.6, 0.6), and ends at the same location as $r_{3}$. Starting from this location, $r_{5}$ ends at (0.4, 0). The weights for segment 151-5 are set as $0.35,0.5,0.85,0.15$ and $1$, respectively. Each segment has two observation locations – one at the start point, another at the middle point.
@@ D.3 Simulation Data Generation Details
D.4 Real-world Data Details
As shown in Figure 5, we construct two subgraphs from PeMS03 dataset. We construct PEMS03-G1 and PEMS03-G2 with temporal distribution shift shown in Table 2. Note that road network structure are required in stream networks. Among all PeMS datasets, PeMS03 is the only one that contains a mapping to real sensor identification number in PeMS. Therefore, other datasets are not available for our setting.
Table 2: Temporal distribution shift of constructed datasets
Dataset Intra-Period Variance Inter-Period Flow Shift
PEMS03-G1 117.4614 21.5453
PEMS03-G2 106.2113 17.4750
(a) PEMS-G1
(b) PEMS-G2
Figure 5: Real-world road network structures and their abstraction for PEMS-G1 and PEMS-G2. In each sub-figure, the left map displays the road network, where freeways are bold gray lines in blue shade, and ramps off the freeway are represented by blue squares. Based on these ramps and road junctions, the network is divided into different segments. Traffic flow monitoring sensors from $\ell_{1}$ to $\ell_{12}$ are placed exclusively on those northbound freeways, marked with green transparent triangles. The right map provides an abstract representation of the road network and sensor locations, using the same symbols for consistency.
D.5 Main Experiment Continued
We evaluated two additional baselines, CopulaTSCP and DeepSTUQ, on the PeMS-G1 dataset, with the results summarized in the Table 3. Using the same backbone model AGCRN, CopulaTSCP significantly underperforms our method and even general CP methods. This is because that learning the joint cumulative distribution function, a requirement for CopulaTSCP, is highly unstable and unreliable in high-dimensional settings like our 12-dimensional joint prediction task. DeepSTUQ, a Bayesian GNN method, provided better efficiency but still failed to outperform STACI. Since DeepSTUQ produces marginal intervals, we aggregated them into a hyper-rectangle for a joint comparison. However, this aggregation lacks a formal guarantee of joint coverage, and applying a formal correction like Bonferroni would result in overly conservative and inefficient prediction sets. These results underscore the limitations of existing methods in high-dimensional joint prediction and highlight the need for specialized approaches.
Table 3: Comparison over additional baselines on the PeMS-G1 dataset.
Method Coverage Efficiency
CopulaCPTS(AGCRN) 70.60 668.00
DeepSTUQ 93.82 66.51
STACI(AGCRN) 95.75 67.54
D.6 Robustness under imperfect topological information
To evaluate our model’s robustness against graph structure noise, we conducted an experiment by perturbing the graphs of the PeMS dataset. A "noisy" graph was constructed by adding 12 spurious edges: for each of the 12 nodes in the original graph, we randomly selected another node and added an edge between them. As shown in the Table 4, the performance of all GNN-based methods degraded when using the noisy graph. Notably, our method, STACI, demonstrated the strongest robustness to these perturbations. It experienced the least performance degradation among all models, showcasing its resilience to imperfect graph structures.
Table 4: Robustness of STACI to noisy sensor locations on the PeMS-G1 and PeMS-G2 datasets across various GNN backbones. Noise Scale represents the standard deviation of the Gaussian noise added to location coordinates. The bolded rows indicate performance on the original, clean data.
Dataset Noise Scale AGCRN ASTGCN STGODE
Coverage Efficiency Coverage Efficiency Coverage Efficiency
G1 0.0 95.75 67.54 $\pm$ 10.94 95.54 67.92 $\pm$ 10.22 95.14 73.62 $\pm$ 9.83
0.5 95.83 71.94 $\pm$ 11.39 95.05 85.24 $\pm$ 14.68 95.09 91.73 $\pm$ 25.11
1.0 95.81 64.12 $\pm$ 10.25 95.05 85.24 $\pm$ 14.68 95.12 99.33 $\pm$ 22.42
2.0 95.58 81.82 $\pm$ 12.73 94.92 90.38 $\pm$ 17.29 95.12 107.26 $\pm$ 24.30
3.0 95.26 119.40 $\pm$ 15.65 95.12 97.44 $\pm$ 14.67 95.12 105.80 $\pm$ 21.06
G2 0.0 95.01 69.62 $\pm$ 12.85 95.05 58.07 $\pm$ 9.83 95.20 74.86 $\pm$ 9.63
0.5 95.01 68.04 $\pm$ 11.06 95.12 69.34 $\pm$ 13.67 95.16 79.08 $\pm$ 10.81
1.0 94.97 69.45 $\pm$ 11.37 95.12 69.34 $\pm$ 13.67 95.14 80.04 $\pm$ 11.29
2.0 95.01 92.31 $\pm$ 13.12 95.03 68.97 $\pm$ 13.18 95.09 79.68 $\pm$ 11.27
3.0 95.09 117.99 $\pm$ 25.89 95.12 69.56 $\pm$ 11.77 95.12 80.06 $\pm$ 11.31
D.7 Running Time
On a single Nvidia Geforce 4080S, the training time of AGCRN with PEMS-G1 is $1419s$. In table 5, we show the computation time of our and baseline CP methods on AGCRN model and PEMS-G1 dataset. Our method STACI has comparable running time with other methods. Here we provide a detailed complexity analysis. Compare to MultiDimSPCI, for each test time point we need: 1) additional estimation of tail-up parameters $\phi,\sigma^{2}$ and using historical covariance matrix weighted addition of spatial covariance and empirical covariance. Consider the node size is $n$ and the optimization methods (e.g., least square in our implementation) iteration round $N$ . The former estimation takes $O(Nn^{2})$, and the latter addition takes $O(n^{3})$ , as pseudo-inverse of matrix is involved. However, MultiDimSPCI also needs matrix inverse for Mahalanobis distance calculation. Additionally, estimation of only two parameters can converge fast. Therefore, our method does not need significantly more time than the baseline method, consistent to the Table 5. STACI provides fast, reliable, and interpretable joint UQ over localized subgraphs across long time horizons.
Table 5: Computation Time (seconds) for Different CP Methods with Different Calibration Set Size $n$
Calibration
Set Size $n$
Sphere
Sphere-ACI
($\gamma=0.01$)
Square MultiDimSPCI
STACI
($\gamma=0$)
STACI
($\gamma=0.01$)
100 $142$ $144$ $24$ $24$ $25$ $25$
200 $152$ $153$ $24$ $24$ $26$ $25$
300 $135$ $129$ $24$ $24$ $25$ $26$
400 $132$ $135$ $25$ $24$ $24$ $25$
500 $135$ $136$ $23$ $24$ $23$ $24$
D.8 Method Details
We provide pseudo-codes for our proposed method in 1. This applies all experiments in this paper, excluding offline setting detailed in Section D.10.
D.9 Hyperparameter Study
The studies for hyperparameter are presented in Figure 6.
(1a)-(1d): AGCRN on PeMS-G1
(2a)-(2d): ASTGCN on PeMS-G1
(3a)-(3d): STGODE on PeMS-G1
(4a)-(4d): AGCRN on PeMS-G2
(5a)-(5d): ASTGCN on PeMS-G2
(6a)-(6d): STGODE on PeMS-G2
Figure 6: Comparison of Coverage and Efficiency for all PeMS experiments with different backbone GNN models and datasets. Each figure show results for different belief weight $\lambda$ and calibration set size $n$, with adaptive step size $\gamma=0$ (a&b) and $0.01$ (c&d). Pre-determined coverage threshold of 9595% is shown by horizontal gray dotted lines.
D.10 Additional Ablation Study: Offline Experiment
@@
In both synthetic and real-world data, MultiDimSPCI achieves the closest to our proposed method, STACI, in efficiency. Therefore, we focus our comparison on four specific variants: vanilla MultiDimSPCI($\gamma=0$), MultiDimSPCI($\gamma=0.01$), STACI($\gamma=0$), and STACI($\gamma=0.01$). In the offline setting, STACI does not update the covariance matrix estimation. To ensure a fair comparison, we similarly fix the covariance matrix for MultiDimSPCI methods at the beginning of the test phase.